# UniRec: Unified Multimodal Encoding for LLM-Based Recommendations

### v1

@@ 2.2 Limitations of Existing Methods
Loss of Schema and Structural Semantics. Text-only LLM-based recommenders (Geng et al., 2022; Li et al., 2023a; Ye et al., 2024) flatten multimodal signals into text prompts. This process erases distinctions between different attribute roles and obscures the nested structure of interactions, making it difficult to preserve schema semantics or accurately model structured data.
Shallow Fusion of Modalities. Conventional multimodal recommenders (He and McAuley, 2016a; Wei et al., 2019b; Tao et al., 2022) typically combine auxiliary signals such as images or reviews through simple concatenation or late fusion. Such shallow integration limits the model’s ability to capture fine-grained cross-modal dependencies or hierarchical structures within user–item interactions.
@@ 3.1 Modality-Specific Encoders
Robust modality-wise encoders and dense, comparable representations are crucial for stable multimodal fusion (Li et al., 2024; Vouitsis et al., 2024; Xu et al., 2022). To this end, we map all inputs into 1024-dimensional embeddings using modality-specific encoders. For text (e.g., titles, reviews), we employ Qwen3-0.6B Embedding [^1] https://huggingface.co/Qwen/Qwen3-Embedding-0.6B. Categorical labels (e.g., product categories) are encoded with the same model using category-aware instructions, avoiding the mixing of sparse one-hot features with dense vectors—a practice known to cause semantic misalignment and training instability (Guan and others, 2022; Li and others, 2022; Cheng and others, 2022). For images, we use CLIP ViT-L/14 (Radford et al., 2021b) [^2] https://huggingface.co/openai/clip-vit-large-patch14, followed by a projection layer that maps the native 768D representations to 1024D. For numerical features, we adopt a Fourier-based Math-Aware Number Encoder (Zhou et al., 2025; Cao et al., 2025), which integrates Fourier components (sine/cosine with log-spaced frequencies), raw magnitude/sign values, and a small learned projection. The encoder is trained with objectives enforcing additivity, invertibility, and distance preservation, while domain-specific adaptations handle temporal cycles (e.g., hour-of-day, month-of-year) and geospatial coordinates projected onto the unit sphere. Implementation details are provided in Appendix A.
@@ 4.1 Experiment Setup
  • Multimodal-Feature Sequential Models. This group adapts classical sequential recommenders by enriching item embeddings with multimodal features. Specifically, GRU4Rec (Hidasi et al., 2016), BERT4Rec (Sun et al., 2019), and SASRec (Kang and McAuley, 2018) are implemented using representations extracted from a pre-trained CLIP ViT-B/32 model (Radford et al., 2021a). CLIP encodes product text and images into a shared embedding space, which replaces ID embeddings while leaving the original model architectures unchanged.
  • Multimodal Recommendation Models. These approaches are explicitly designed to integrate multimodal signals into recommendation. VBPR (He and McAuley, 2016a) augments matrix factorization with visual preference vectors derived from product images. MMGCN (Wei et al., 2019b) builds modality-specific user–item graphs and aggregates them with graph neural networks. BM3 (Zhou et al., 2023b) introduces a self-supervised learning framework that applies dropout to generate contrastive views without explicit negative sampling, improving robustness. LGMRec (Guo et al., 2024) decouples collaborative filtering and content modeling by combining local modality-specific graphs with global hypergraph-based embeddings, achieving stronger alignment between content and interaction signals.
  • LLM-based Multimodal Recommenders. Recent methods leverage large language models to reason over multimodal content. IISAN (Fu et al., 2024) adopts a parameter-efficient fine-tuning framework that decouples intra- and inter-modal adaptation, improving both training efficiency and scalability while maintaining accuracy. MLLM-MSR (Ye and others, 2025) takes a summarization approach: it converts product images into textual descriptions, uses an LLM to summarize user histories, and fine-tunes the model for sequential recommendation. These models highlight the emerging trend of directly coupling multimodal inputs with LLM.
@@ 4.3 Ablation Studies Validate the Design of UniRec
  • CLIP-Style Projection. Maps each modality into a shared latent space through modality-specific linear layers, with alignment guided by a contrastive objective, following dual-encoder vision–language paradigms such as CLIP and ALIGN (Radford et al., 2021b; Jia et al., 2021).
@@ 5 Related Work
Multimodal Recommendation. A large body of research has shown that incorporating auxiliary modalities such as text, images, or reviews can substantially enrich user and item representations and alleviate data sparsity (He and McAuley, 2016b; Wei et al., 2019a; Tao and others, 2022; Zhou et al., 2023a; Liu and others, 2023; Yu et al., 2025). These works established that multimodal signals carry complementary semantics beyond ID-based interactions and can improve personalization in various domains. More recent efforts have begun to align multimodal content with pretrained language models, as in VIP5 (Geng et al., 2023), demonstrating the promise of combining vision, language, and recommendation. Collectively, these advances highlight multimodality as a powerful driver for next-generation recommender systems.
@@ Appendix A Encoder Implementation Details
Images are processed with CLIP ViT-L/14 (Radford et al., 2021b), which produces 512-dimensional embeddings aligned with text through large-scale contrastive pretraining. This backbone offers a strong balance between representational quality and computational efficiency, enabling reliable multimodal alignment.

### v2

@@ 2.2 Limitations of Existing Methods
Loss of Schema and Structural Semantics. Text-only LLM-based recommenders (Geng et al., 2022; Li et al., 2023a; Ye et al., 2025) flatten multimodal signals into text prompts. This process erases distinctions between different attribute roles and obscures the nested structure of interactions, making it difficult to preserve schema semantics or accurately model structured data.
Shallow Fusion of Modalities. Conventional multimodal recommenders (He and McAuley, 2016; Wei et al., 2019b; Tao et al., 2022) typically combine auxiliary signals such as images or reviews through simple concatenation or late fusion. Such shallow integration limits the model’s ability to capture fine-grained cross-modal dependencies or hierarchical structures within user–item interactions.
@@ 3.1 Modality-Specific Encoders
Robust modality-wise encoders and dense, comparable representations are crucial for stable multimodal fusion (Li et al., 2024; Vouitsis et al., 2024; Xu et al., 2022). To this end, we map all inputs into 1024-dimensional embeddings using modality-specific encoders. For text (e.g., titles, reviews), we employ Qwen3-0.6B Embedding [^1] https://huggingface.co/Qwen/Qwen3-Embedding-0.6B. Categorical labels (e.g., product categories) are encoded with the same model using category-aware instructions, avoiding the mixing of sparse one-hot features with dense vectors—a practice known to cause semantic misalignment and training instability (Guan and others, 2022; Li and others, 2022; Cheng and others, 2022). For images, we use CLIP ViT-L/14 (Radford et al., 2021) [^2] https://huggingface.co/openai/clip-vit-large-patch14, followed by a projection layer that maps the native 768D representations to 1024D. For numerical features, we adopt a Fourier-based Math-Aware Number Encoder (Zhou et al., 2025; Cao et al., 2025), which integrates Fourier components (sine/cosine with log-spaced frequencies), raw magnitude/sign values, and a small learned projection. The encoder is trained with objectives enforcing additivity, invertibility, and distance preservation, while domain-specific adaptations handle temporal cycles (e.g., hour-of-day, month-of-year) and geospatial coordinates projected onto the unit sphere. Implementation details are provided in Appendix A.
@@ 4.1 Experiment Setup
  • Multimodal-Feature Sequential Models. This group adapts classical sequential recommenders by enriching item embeddings with multimodal features. Specifically, GRU4Rec (Hidasi et al., 2016), BERT4Rec (Sun et al., 2019), and SASRec (Kang and McAuley, 2018) are implemented using representations extracted from a pre-trained CLIP ViT-B/32 model (Radford et al., 2021). CLIP encodes product text and images into a shared embedding space, which replaces ID embeddings while leaving the original model architectures unchanged.
  • Multimodal Recommendation Models. These approaches are explicitly designed to integrate multimodal signals into recommendation. VBPR (He and McAuley, 2016) augments matrix factorization with visual preference vectors derived from product images. MMGCN (Wei et al., 2019b) builds modality-specific user–item graphs and aggregates them with graph neural networks. BM3 (Zhou et al., 2023b) introduces a self-supervised learning framework that applies dropout to generate contrastive views without explicit negative sampling, improving robustness. LGMRec (Guo et al., 2024) decouples collaborative filtering and content modeling by combining local modality-specific graphs with global hypergraph-based embeddings, achieving stronger alignment between content and interaction signals.
  • LLM-based Multimodal Recommenders. Recent methods leverage large language models to reason over multimodal content. IISAN (Fu et al., 2024) adopts a parameter-efficient fine-tuning framework that decouples intra- and inter-modal adaptation, improving both training efficiency and scalability while maintaining accuracy. MLLM-MSR (Ye et al., 2025) takes a summarization approach: it converts product images into textual descriptions, uses an LLM to summarize user histories, and fine-tunes the model for sequential recommendation. These models highlight the emerging trend of directly coupling multimodal inputs with LLM.
@@ 4.3 Ablation Studies Validate the Design of UniRec
  • CLIP-Style Projection. Maps each modality into a shared latent space through modality-specific linear layers, with alignment guided by a contrastive objective, following dual-encoder vision–language paradigms such as CLIP and ALIGN (Radford et al., 2021; Jia et al., 2021).
@@ 5 Related Work
Multimodal Recommendation. A large body of research has shown that incorporating auxiliary modalities such as text, images, or reviews can substantially enrich user and item representations and alleviate data sparsity (He and McAuley, 2016; Wei et al., 2019a; Tao and others, 2022; Zhou et al., 2023a; Liu and others, 2023; Yu et al., 2025). These works established that multimodal signals carry complementary semantics beyond ID-based interactions and can improve personalization in various domains. More recent efforts have begun to align multimodal content with pretrained language models, as in VIP5 (Geng et al., 2023), demonstrating the promise of combining vision, language, and recommendation. Collectively, these advances highlight multimodality as a powerful driver for next-generation recommender systems.
@@ Appendix A Encoder Implementation Details
Images are processed with CLIP ViT-L/14 (Radford et al., 2021), which produces 512-dimensional embeddings aligned with text through large-scale contrastive pretraining. This backbone offers a strong balance between representational quality and computational efficiency, enabling reliable multimodal alignment.